OpenStudy (anonymous):

Suppose a building inspector orders the height of a ceiling to be 8 feet tall, but the workers buy lumber that is only 50 centimeters tall. How many pieces of lumber will they have to stack to create an 8-foot ceiling? Explain your answer. (Hint: 1 inch = 2.54 centimeters.)

OpenStudy (anonymous):

12inches=1feet

OpenStudy (anonymous):

2.54cm=1inch 50cm=x x=50/2.54=19.6850394inhes

OpenStudy (anonymous):

8 foot=12*8inches=96inhes

OpenStudy (anonymous):

therfoe no. of lumbers=96/19.68=4.87=5(appx)